OpenStudy (anonymous):

Two cones have a scale factor of 2 : 5. The smaller cone has a surface area of 96(pi) square yards. Find the surface area of the larger cone in terms of pi. Help if you can.

OpenStudy (anonymous):

area goes with the square, so if the scale factor is 2:5 area ratio is 4 : 25

OpenStudy (anonymous):

How would I solve this? I am completely lost. Not to mention but my entire class also had wrong answers.

OpenStudy (anonymous):

\[96\pi\] is the smaller one right?

OpenStudy (anonymous):

you want the larger one solve \[\frac{25}{4}=\frac{x}{96\pi}\] via \[x=\frac{25\times 96\pi}{4}\] canecel first multiply last

OpenStudy (anonymous):

What would I cancel out? I don't see anything to cancel out. If nothing, I come to 600(pi)

OpenStudy (anonymous):

Which sounds incredibly wrong.

OpenStudy (anonymous):

4 goes in to 96 evenly

OpenStudy (anonymous):

But it can't go evenly into 25. Wouldn't I need to do it to both?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\frac{6\times 2}{3}=2\times 2=4\] 3 goes in to 6 but not in to 2

OpenStudy (anonymous):

in any case your answer \(600\pi\) is right

OpenStudy (anonymous):

Alright then thank you.

OpenStudy (anonymous):

not sure why you think it is incredibly wrong

OpenStudy (anonymous):

we can do another easy example if you like

OpenStudy (anonymous):

I have another problem along the same lines if you want to do that as well.

OpenStudy (anonymous):

ok if i can

OpenStudy (anonymous):

Two spheres have a scale factor of 3 : 8. The smaller sphere has a volume of 54(pi) cubic meters. Find the volume of the larger sphere in terms of Pi.

OpenStudy (anonymous):

area goes with the square, volume with the cube like cubic inches etc

OpenStudy (anonymous):

what is 3 cubed?

OpenStudy (anonymous):

i.e. \[3^3\] or \[3\times 3\times 3\]?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

good and 8 cubed ?

OpenStudy (anonymous):

512

OpenStudy (anonymous):

solve \[\frac{512}{27}=\frac{x}{54\pi}\]

OpenStudy (anonymous):

via \[x=\frac{512\times 54\pi}{27}\] cancel first multiply last

OpenStudy (anonymous):

1024(pi)?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Thanks!